OpenStudy (anonymous):

example from book that i don't fully understand:

OpenStudy (anonymous):

\[\log_{10} 2^{1.42}\]

OpenStudy (anonymous):

according to this a theorem: this is equal to; 1.42log2 =(1.42)(0.30103) 2^1.42=2.6758 how did they get the last part?

OpenStudy (experimentx):

calculator ...

OpenStudy (anonymous):

i can't use a calculator in the exam..

OpenStudy (experimentx):

do you get log tables??

OpenStudy (anonymous):

the book said use theorem 7 to compute \[2^{1.42}\] and theorem 7 is: \[ \log_{a} X ^{y} = y \log_{a} x\]

OpenStudy (anonymous):

nope only some logs like log2 and log3..

OpenStudy (anonymous):

@lgbasallote

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=log+base+10+2^1.42 something fishy

OpenStudy (anonymous):

i don't know but the answer for 2^1.42 = 2.6758

OpenStudy (lgbasallote):

that's according to a calculator

OpenStudy (anonymous):

we just need to solve it using the theorem.

OpenStudy (lgbasallote):

it's not the answer to \[\log_{10} 2^{1.42}\]it's just telling you the value of \[2^{1.42}\]

OpenStudy (lgbasallote):

maybe to compare

OpenStudy (anonymous):

i know that \[\log_{10} 2^{1.42} = 0.42746\] but we need to find 2^1.42

OpenStudy (anonymous):

using logs

OpenStudy (lgbasallote):

dint you just say \[2^{1.42} = 2.6758\]

OpenStudy (lgbasallote):

anyway..you cant solve \[2^{1.42}\]without using a calculator...unless you know calculus or some other high level math

OpenStudy (anonymous):

okay, i guess then i can't do my exercise.. thank you for all the help.

OpenStudy (anonymous):

@waterineyes @satellite73

OpenStudy (anonymous):

Which value you want to find : \(2^{1.42}\) Or: \(log(2^{1.42})\) ??

OpenStudy (anonymous):

the first using the second

OpenStudy (anonymous):

And you are not allowed to use calculators ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then I think you are allowed to Log Tables in exams...

OpenStudy (anonymous):

maybe the book is wrong because i have said everything it has told me

OpenStudy (anonymous):

Can you use Log Tables in exams or not ??

OpenStudy (anonymous):

no log table but we are given certain logs that are needed

OpenStudy (anonymous):

Can we use approximation here ??

OpenStudy (anonymous):

i guess..

OpenStudy (experimentx):